OpenStudy (anonymous):

The sum of 2 numbers is 35. twice the first number is 25 more than half of the second numbe. Find the numbers

OpenStudy (anonymous):

let x the 1st no. y the 2nd no. so if you'll analyze the question: "The sum of 2 numbers is 35" that would mean : x+y = 35 do you agree?

OpenStudy (anonymous):

how about the 2nd one??

OpenStudy (anonymous):

can you figure out what "twice the first number" means?

OpenStudy (anonymous):

No :c

OpenStudy (anonymous):

it means 2x do you agree?

OpenStudy (anonymous):

r you typing?

OpenStudy (anonymous):

please give me the equation :(

OpenStudy (anonymous):

i will but do you understand that?

OpenStudy (anonymous):

yes!!

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

OMG <33 thanks!!! I really need it

OpenStudy (anonymous):

2x = 25 + y/2 now you can substitute

OpenStudy (anonymous):

do you understand? remember that the 1st eq is : x+y = 35 you can solve for x: x=35-y

OpenStudy (anonymous):

can it be by elimation?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

how to?

OpenStudy (anonymous):

substitution is easier :) since x = 35-y